# On Positional Bias of Faithfulness for Long-form Summarization

### v1

@@ Abstract
Large Language Models (LLMs) often exhibit positional bias in long-context settings, under-attending to information in the middle of inputs. We investigate the presence of this bias in long-form summarization, its impact on faithfulness, and various techniques to mitigate this bias. To consistently evaluate faithfulness, we first compile a benchmark of eight human-annotated long-form summarization datasets and perform a meta-evaluation of faithfulness metrics. We show that LLM-based faithfulness metrics, though effective with full-context inputs, remain sensitive to document order, indicating positional bias. Analyzing LLM-generated summaries across six datasets, we find a "U-shaped" trend in faithfulness, where LLMs faithfully summarize the beginning and end of documents but neglect middle content. Perturbing document order similarly reveals models are less faithful when important documents are placed in the middle of the input. We find that this behavior is partly due to shifting focus with context length: as context increases, summaries become less faithful, but beyond a certain length, faithfulness improves as the model focuses on the end. Finally, we experiment with different generation techniques to reduce positional bias and find that prompting techniques effectively direct model attention to specific positions, whereas more sophisticated approaches offer limited improvements. Our data and code are available in https://github.com/meetdavidwan/longformfact.
@@
- Simply prompting the models to focus on the different locations of the documents is effective.
@@ 1 Introduction
Finally, we investigate methods that attempt to mitigate such positional bias. We explore methods such as prompting to focus on certain parts, hierarchical merging, incremental updating, and calibration methods. We find that prompting methods are effective at improving summary’s faithfulness towards certain positions, while more sophisticated methods struggle to address this issue.
@@ 2.1 Long-form Summarization
We consider the task of summarization, where a model generates an $m$-sentence summary $S=\{s_{1},s_{2},...,s_{m}\}$ from input document(s) $D$. For long-form summarization tasks, such as multi-document summarization, the input $D$ consists of $n$ documents: $D=\{d_{1},d_{2},...,d_{n}\}$. We refer to the boundaries of these documents as natural document boundaries, since there are no restrictions on how long each document may be. To unify different datasets, we can similarly split a single-document dataset, such as a scientific document, into different sections and refer to these sections as “documents.” Alternatively, one may split the document into fixed-length chunks of words or tokens.
@@ 3.1 LongFormFact
is a multi-document news summarization dataset that focuses on summarizing diverse or conflicting information. The authors collected sentence-level binary faithfulness judgments.
Metric Doc Merge MN QM GR PB AX SQ CS DS Average
MiniCheck Original 52.5 62.8 60.8 72.2 56.1 77.5 49.5 48.6 60.0
Min 50.1 62.8 60.8 50.8 50.0 77.5 54.2 49.6 57.0
Mean 54.8 62.8 60.8 51.5 55.5 77.5 54.2 49.4 58.3
Max 54.7 62.8 60.8 73.1 62.3 77.5 56.9 43.5 61.4
GPT-4o Full 52.2 64.3 61.6 80.1 57.6 76.7 63.4 57.4 64.1
Min 50.0 60.7 60.0 54.4 52.1 80.6 54.2 50.3 57.8
Mean 46.3 60.7 60.0 64.3 56.2 80.6 54.2 56.6 59.9
Max 60.0 60.7 60.0 76.0 55.6 80.6 60.6 53.0 63.3
@@ 3.2 Experimental Setup
We primarily experiment using GPT-4o as the backbone LLM for the metric. We also explore the applicability of MiniCheck Tang et al. (2024b) – an automatic metric that has demonstrated efficacy comparable to GPT-4 performance – to the long-context setting. We note that MiniCheck by default splits the input into chunks of fixed number of tokens and takes the maximum over the documents, which we include as one of the baselines. Additionally, we present results using the Llama-3.1-8B model in Appendix B.2.
@@
3.3 Results
The main results are shown in Table 3. For MiniCheck, we observe that taking the maximum over the document achieves the highest correlations on average. This shows that this document aggregation method achieves the best results even for long-form summarization. Interestingly, the original strategy of taking the maximum over fixed input context performs on average $1.4\%$ lower than using natural document boundaries, suggesting that only evaluating the relevant context is important.
@@ 3.3 Results
In addition to standard meta-evaluation, we also perform an analysis by ordering the documents in terms of importance. Specifically, for datasets where document boundaries exist, we calculate the importance of each document relative to the model-generated summary using sentence similarity. [^1] We use SentenceTransformer Reimers and Gurevych (2019) with the all-mpnet-base-v2 model. We then order the documents into top (beginning), middle, and bottom (end), corresponding to the placement of the most important documents. Note that the only change from the regular case is the document order, which should have no effect for MultiNews and DiverseSumm, where the documents are in random order, but may have an effect for the other datasets where it breaks the natural flow of the document. The results are presented in Table 4. In addition to BACC, we also include sensitivity, which represents the difference between the largest and lowest scores across each dataset and each document order.
Overall, we find that the metric is sensitive to document order, with high sensitivity observed across each dataset and reaching a 7.8% difference for MultiNews. When comparing only the top, middle, and bottom orderings, we find no clear trend across datasets; however, on average, the sensitivity is highest when the important document is placed in the middle. This indicates that the metric achieves the lowest balanced accuracy (BACC) when the important document is in the middle, whereas placing it at the top results in the smallest difference, suggesting that the LLM has a stronger lead bias, i.e. performing better when the important documents are at the beginning of the input. Therefore, it may still be beneficial to use the metric that evaluates each document individually and aggregates the results via the maximum operation to reduce positional bias. This approach achieves a similar BACC compared to the full-context setting while inherently not being sensitive to input order.
@@ 4.1 Experimental Setup
We include two representative multi-document summarization datasets, MultiNews Fabbri et al. (2019) and MultiXScience Lu et al. (2020); two long-form summarization datasets, ArXiv and PubMed Cohan et al. (2018); and two recent summarization datasets with extremely long contexts, DiverseSumm Huang et al. (2024) and SummHay Laban et al. (2024). For ArXiv, PubMed, MultiNews, and MultiXScience, we randomly sample 100 examples from the validation set, each consisting of five documents or sections. For DiverseSumm, we use all original 10 documents and randomly sample 100 examples.
@@
Though we observe that the faithfulness metric using GPT-4o achieves better performance in Section 3.3, we use MiniCheck due to efficiency and cost considerations. In Appendix C.3, we show that MiniCheck exhibits similar trends as the faithfulness with GPT-4o.
@@ 4.2 Faithfulness Analysis
Our main analysis evaluates whether the summary is more faithful to documents in certain positions. To do so, we calculate the Minicheck faithfulness score with respect to all documents individually. We note that if we directly report the summary’s faithfulness to each document, it may not accurately reflect the actual faithfulness, as it is also confounded by coverage, i.e., how much the summary draws on content from each document. In fact, Huang et al. (2024) use the faithfulness score per document to measure coverage bias, and we provide an analysis of this coverage in Appendix C.2. To remove the effect of coverage, we consider attribution; that is, determining which document each sentence of the summary is discussing. For each summary sentence, we take the maximum faithfulness score over all documents. This is, in fact, the same process as one of the best document merging strategies reported in Section 4.4.
@@
ArXiv 75.2 77.2 73.7 71.8 5.4
PubMed 82.3 80.6 83.5 80.4 3.1
MultiNews 83.1 83.5 81.0 83.1 2.5
MultiXScience 76.0 75.6 73.5 75.6 2.5
SummHay 75.6 66.4 68.7 85.5 19.1
Sensitivity - 2.7 4.5 5.1 1.8 -
@@
Results.
We present the results in Figure 2. Generally, we observe that the models exhibit a dominant U-shaped curve, particularly on ArXiv, PubMed, and MultiNews. On MultiXScience, we observe a strong lead bias, another prevalent trend in abstractive summarization. The trend for DiverseSumm is interesting, as it is more of a linear trend. This is partly due to the nature of the task, where we explicitly ask to perform synthesis across the documents, which helps the model to focus on different parts. On SummHay, while the three models behave differently, we note that they all achieve lower faithfulness scores for the middle document.
@@ 4.3 Perturbed Input Analysis
Next, we conduct a similar perturbation analysis as described in Section 3.3, where we reorder the documents according to importance. Here, we determine the importance using the similarity between each document and the reference summary. We exclude DiverseSumm, as it does not contain reference summaries. The results are reported in Table 5. We observe a similar trend when we analyze the perturbation for the metric: The models generate the most faithful summaries either with the original order or when the important document is placed at the front. When looking at the sensitivity across each ordering, the middle case has the highest sensitivity, indicating that the summaries are least faithful in this scenario. Interestingly, we observe the lowest sensitivity for the bottom case, but this is due to the large variation in SummHay, where the bottom ordering achieves a score 19.1 points over the top case, representing the least faithful summary for this dataset. We posit that this is because of how models handle long contexts, i.e., focusing towards the end when the context increases, which we confirm in the subsequent analysis on the correlation between increasing context length and faithfulness.
@@ 4.4 Faithfulness and Length Correlation
Results are in Figure 3. For ArXiv, PubMed, and DiverseSumm, we observe the same U-shape occurring. This indicates that the faithfulness of the summary decreases as more documents are included up to a certain point, after which it begins to improve. This suggests that the model may switch modes at a certain context length and focus more on the documents introduced later. On MultiNews, we can similarly observe this trend, as the faithfulness steadily improves for most models as we include more input documents. This phenomenon is consistent with prior research on LLMs’ behavior of summarizing extremely long input, where models focus a large portion of their attention on the later part Kim et al. (2024b); Laban et al. (2024).
@@ 5.2 Results
We present the results of running the different methods with GPT-4o in Figure 4. At the top of the figure, the prompting methods are displayed. Generally, instructing the model to focus on different positions is effective and improves the faithfulness for that specific position. For example, focusing on the bottom achieves higher faithfulness for the last document than the original prompt for 4 out of 5 datasets. Similarly, the “focus middle” prompt achieves higher faithfulness than the original case for the middle document in all cases except for DiverseSumm, where the method improves faithfulness for the 6th document instead of the 5th one.
Interestingly, the more sophisticated methods illustrated at the bottom of Figure 4 perform worse than the original prompt across all datasets. The only exception is the incremental approach on MultiXScience. It is noteworthy, as we generally observe a very different trend compared to the other summarization techniques. For example, on ArXiv, PubMed, and DiverseSumm, incremental updating fails to generate summaries for the middle documents, as indicated by gaps in the corresponding lines on the plot. This is primarily because the method keeps only a working summary cache and asks the model to update it based on the next document, leading to a strong recency bias: the model either retains its current summary (i.e., remains faithful to the first document) or updates its summary based on the last document. Thus, while this method seems to improve faithfulness, it essentially transfers from one positional bias to another—that is, from "lost-in-the-middle" to recency bias.
@@ 6 Related Work
To improve the benchmarking of summarization systems, many studies collect human annotations of faithfulness judgments to create meta-evaluation benchmarks that measure the effectiveness of faithfulness metrics Fabbri et al. (2021); Pu et al. (2023); Tang et al. (2023); Goyal et al. (2023); Zhang et al. (2024c); Liu et al. (2023b, 2024b). This effort has led to the development of strong faithfulness metrics Laban et al. (2022); Fabbri et al. (2022); Zha et al. (2023). More recently, LLM-based metrics, which prompt powerful LLMs to perform evaluations Liu et al. (2023a); Wang et al. (2023); Fu et al. (2024), have shown high correlations with human judgments. Subsequently, MiniCheck Tang et al. (2024b) has focused on distilling such knowledge into smaller NLI models, combining both strong performance and efficiency. While most of these studies concentrate on standard summarization tasks, few address long-form summarization Krishna et al. (2023); Zhang et al. (2024b); Huang et al. (2024). Specifically, Zhang et al. (2024b) collate a meta-evaluation benchmark and develop an automatic metric for a large collection of long-form summarization. We further extend their work by incorporating additional benchmarks, analyzing the performance of the latest faithfulness evaluation metrics, and, more importantly, investigating their sensitivity to positional bias.

### v2

@@ Abstract
Large Language Models (LLMs) often exhibit positional bias in long-context settings, under-attending to information in the middle of inputs. We investigate the presence of this bias in long-form summarization, its impact on faithfulness, and various techniques to mitigate this bias. To consistently evaluate faithfulness, we first compile a benchmark of eight human-annotated long-form summarization datasets and perform a meta-evaluation of faithfulness metrics. We show that LLM-based faithfulness metrics, though effective with full-context inputs, remain sensitive to document order, indicating positional bias. Analyzing LLM-generated summaries across six datasets, we find a "U-shaped" trend in faithfulness, where LLMs faithfully summarize the beginning and end of documents but neglect middle content. Perturbing document order similarly reveals models are less faithful when important documents are placed in the middle of the input. We find that this behavior is partly due to shifting focus with context length: as context increases, summaries become less faithful, but beyond a certain length, faithfulness improves as the model focuses on the end. Finally, we experiment with different generation techniques to reduce positional bias and find that prompting techniques direct model attention to specific positions, whereas more sophisticated approaches offer limited improvements. Our data and code are available in https://github.com/meetdavidwan/longformfact.
@@
- Simply prompting the models to focus on the different locations of the documents is moderately effective.
@@ 1 Introduction
Finally, we investigate methods that attempt to mitigate positional bias. We explore methods such as prompting to focus on certain parts, hierarchical merging, incremental updating, and calibration methods. We find that prompting methods are moderately effective at improving summary’s faithfulness towards certain positions, while more sophisticated methods struggle to address this issue.
@@ 2.1 Long-form Summarization
We consider the task of summarization, where a model generates an $m$-sentence summary $S=\{s_{1},s_{2},...,s_{m}\}$ from input document(s) $D$. For long-form summarization tasks, we follow Ravaut et al. (2024) and generally consider that the documents need to contain at least 2k tokens. For this task, such as multi-document summarization, the input $D$ consists of $n$ documents: $D=\{d_{1},d_{2},...,d_{n}\}$. We refer to the boundaries of these documents as natural document boundaries, since there are no restrictions on how long each document may be. To unify different datasets, we can similarly split a single-document dataset, such as a scientific document, into different sections and refer to these sections as “documents.” Alternatively, one may split the document into fixed-length chunks of words or tokens. To generate summaries, the entire input is truncated to fit the context window of each respective model.
@@ 3 Faithfulness Metrics Meta-Evaluation
Metric Doc Merge MN QM GR PB AX SQ CS DS Average
MiniCheck Original 53.8 53.9 50.0 79.7 50.0 83.0 50.9 55.2 59.6
Min 54.7 55.7 55.1 50.0 50.0 83.0 54.2 49.6 56.5
Mean 64.5 55.7 55.1 52.6 53.4 83.0 54.2 46.0 58.1
Max 56.6 55.7 55.1 84.8 62.3 83.0 59.7 49.3 63.3
GPT-4o Full 52.2 64.3 61.6 80.1 57.6 76.7 63.4 57.4 64.1
Min 50.0 60.7 60.0 54.4 52.1 80.6 54.2 50.3 57.8
Mean 46.3 60.7 60.0 64.3 56.2 80.6 54.2 56.6 59.9
Max 60.0 60.7 60.0 76.0 55.6 80.6 60.6 53.0 63.3
@@ 3.1 LongFormFact
is a multi-document news summarization dataset that focuses on conflicting information. The authors collected sentence-level binary faithfulness judgments.
@@ 3.2 Experimental Setup
We primarily experiment using GPT-4o as the backbone LLM for the metric. We also explore the applicability of MiniCheck [^1] We use Bespoke-MiniCheck-7B. Tang et al. (2024b) – an automatic metric that has demonstrated efficacy comparable to GPT-4 performance – to the long-context setting. We note that MiniCheck by default splits the input into chunks of fixed number of tokens and takes the maximum over the documents, which we include as one of the baselines. Additionally, we present results using the Llama-3.1-8B model in Appendix B.2.
@@
3.3 Results
The main results are shown in Table 3. For MiniCheck, we observe that taking the maximum over the document achieves the highest correlations on average. This shows that this document aggregation method achieves the best results even for long-form summarization. Interestingly, the original strategy of taking the maximum over fixed input context performs on average $3.7\%$ lower than using natural document boundaries, suggesting that only evaluating the relevant context is important. We also note that MiniCheck trails the strongest GPT-4o-based metric by only $0.8\%$, while matching the performance of GPT-4o-based metric with the same aggregation method.
@@ 3.3 Results
In addition to standard meta-evaluation, we also perform an analysis by ordering the documents in terms of importance. Specifically, for datasets where document boundaries exist, we calculate the importance of each document relative to the model-generated summary using sentence similarity. [^2] We use SentenceTransformer Reimers and Gurevych (2019) with the all-mpnet-base-v2 model. We then order the documents into top (beginning), middle, and bottom (end), corresponding to the placement of the most important documents. To illustrate, assume there are five documents with importance ranks of 1, 3, 2, 5, and 4, where rank 1 denotes the most important and rank 5 the least. The "top" ordering would sort documents by importance (e.g., 1-2-3-4-5), the "bottom" ordering would prioritize the least important documents (e.g., 5-3-4-2-1), and the "middle" ordering would reflect a mid-tier arrangement (e.g., 4-2-1-3-5). Note that the only change from the regular case is the document order, which should have no effect for MultiNews and DiverseSumm, where the documents are in random order, but may have an effect for the other datasets where it breaks the natural flow of the document.The results are presented in Table 4. In addition to BACC, we also include sensitivity, defined as the maximum difference between scores computed using the original ordering and those with different orderings.
Overall, we find that the metric is sensitive to document order, with high sensitivity observed across each dataset and reaching a 7.8% difference for MultiNews. When comparing only the top, middle, and bottom orderings, we find no clear trend across datasets; however, on average, the sensitivity is highest when the important document is placed in the middle. This indicates that the metric achieves the lowest BACC when the important document is in the middle, whereas placing it at the top results in the smallest difference, suggesting that the LLM has a stronger lead bias, i.e. performing better when the important documents are at the beginning of the input. Therefore, it may still be beneficial to use the metric that evaluates each document individually and aggregates the results via the maximum operation to reduce positional bias. This approach achieves a similar BACC compared to the full-context setting while inherently not being sensitive to input order.
@@
Tasks Doc Split Doc Words Summ Split Summ Words
MultiXScience 5 804.9 6.9 186.0
PubMed 5 2850.3 7.4 190.1
MultiNews 5 4925.5 8.5 215.3
ArXiv 5 5825.5 7.1 181.6
DiverseSumm 10 7561.5 16.0 452.5
SummHay 100 87913.1 7.5 52.2
Table 5: Statistics for the generated summaries.
@@ 4.1 Experimental Setup
We include two representative multi-document summarization datasets, MultiNews Fabbri et al. (2019) and MultiXScience Lu et al. (2020); two long-form summarization datasets, ArXiv and PubMed Cohan et al. (2018); and two recent summarization datasets with extremely long contexts, DiverseSumm Huang et al. (2024) and SummHay Laban et al. (2024). For ArXiv, PubMed, MultiNews, and MultiXScience, we randomly sample 100 examples from the validation set, each consisting of five documents or sections. For DiverseSumm, we use all original 10 documents and randomly sample 100 examples. The dataset statistics are shown in Table 5.
@@
Although GPT-4o demonstrates better faithfulness performance in Section 3.3, we choose MiniCheck for its nearly comparable performance, as well as its greater efficiency and lower cost. In Appendix C.6, we show evaluating with GPT-4o-based metric on a small subset, which exihibits similar trends.
@@ 4.2 Faithfulness Analysis
Our main analysis evaluates whether the summary is more faithful to documents in certain positions. To do so, we calculate the Minicheck faithfulness score with respect to all documents individually. We note that if we directly report the summary’s faithfulness of each document, it may not accurately reflect faithfulness, as it is also confounded by coverage, i.e., how much the summary draws on content from each document. In fact, Huang et al. (2024) use the faithfulness score per document to measure coverage bias, and we provide an analysis of this coverage in Appendix C.5 as well as an analysis of document content overlap in Appendix C.1. To remove the effect of coverage, we consider attribution; that is, determining which document each sentence of the summary is discussing. For each summary sentence, we take the maximum faithfulness score over all documents. This is, in fact, the same process as one of the best document merging strategies reported in Section 4.4.
@@
Results.
We present the results in Figure 2. Generally, we observe that the models exhibit a dominant U-shaped curve, particularly on ArXiv, PubMed, MultiNews, and MultiXScience, where the middle documents are less faithful than the first or last documents. While most models exhibit this trend, Mixtral shows a more pronounced lead bias, especially on ArXiv. The trend for DiverseSumm is interesting, as it is more of a linear trend. This is partly due to the nature of the task, where we explicitly ask to perform synthesis across the documents, which helps the model to focus on different parts. On SummHay, although the three models behave differently, we observe that GPT-4o and Claude yield lower faithfulness scores for the middle document. Interestingly, Gemini is the only model that exhibits the reverse trend, performing better on the middle documents; however, its average faithfulness score is not as high as that of the other two models. Appendix C.3 includes further discussions on variance across tasks.
@@ Results.
ArXiv 91.3 91.5 92.2 91.7 0.9
PubMed 91.3 92.6 93.4 92.2 2.1
MultiNews 92.2 90.0 89.0 90.6 3.2
MultiXScience 70.1 67.6 62.1 68.2 8.0
SummHay 73.5 62.7 66.5 84.6 11.1
Avg. Sensitivity - - 3.4 4.3 3.2 -
@@ 4.3 Perturbed Input Analysis
Next, we conduct a similar perturbation analysis with GPT-4o as described in Section 3.3, where we reorder the documents according to importance. Here, importance is determined using the similarity between each document and the reference summary. We exclude DiverseSumm, as it does not contain reference summaries. The results are reported in Table 6. We observe a similar trend when we analyze the perturbation for the metric: The models generally generate the most faithful summaries either with the original order or when the important document is placed at the front. When looking at the sensitivity across each ordering, the middle case has the highest sensitivity. Interestingly, we observe the bottom ordering achieves a score 21.9 points over the top case for SummHay. We posit that this is because of how models handle long contexts, i.e., focusing towards the end when the context increases, which we confirm in the subsequent analysis on the correlation between increasing context length and faithfulness.
@@ 4.4 Faithfulness and Length Correlation
Results are in Figure 3. In MultiXScience, which contains relatively few words per document, we observe a strong lead bias. However, when moving to long-form summarization on datasets such as Pubmed and ArXiv, we observe a U-shaped trend: summary faithfulness decreases as the number of documents grows, then begins to improve beyond a certain threshold. This suggests that the model may “switch modes” at a particular context length and focus more on documents introduced later. On MultiNews, a similar pattern appears, as faithfulness steadily increases for most models with additional input documents. These observations are consistent with prior research on LLMs’ behavior with extremely long inputs, where models concentrate much of their attention on the later sections Kim et al. (2024b); Laban et al. (2024).
@@ 5.2 Results
We present the results of running the different methods with GPT-4o in Figure 4. The top portion of the figure shows the prompting-based methods. Instructing the model to focus on either the top or bottom documents proves effective for improving faithfulness at those specific positions. For instance, the “focus top” prompt yields higher faithfulness scores for the last document than the original prompt in 4 out of 5 datasets. The “focus middle” prompt achieves higher faithfulness than the original case for the middle document on ArXiv and DiverseSumm. However, we observe that the model often fails to follow the instructions adequately to focus on the middle documents: “focus top” achieves the best faithfulness for the middle documents on PubMed, while “focus bottom” performs best on DiverseSumm. This suggests that prompting can partially alleviate positional bias.
Interestingly, the more sophisticated methods illustrated at the bottom of Figure 4 perform worse than the original prompt across all datasets. It is noteworthy, as we generally observe a very different trend for methods that change the summarization protocol. For example, on ArXiv, PubMed, and DiverseSumm, incremental updating contains gaps in the corresponding lines on the plot. This issue arises because the method maintains only a working summary cache and updates it solely based on the next document, leading to a pronounced recency bias: the model either retains its current summary and thus remains faithful to the first document or focuses on the last document. While this approach may appear to improve faithfulness, it ultimately replaces one positional bias, i.e., lost-in-the-middle, with another, i.e., recency bias.
@@ 6 Related Work
To improve the benchmarking of summarization systems, many studies collect human annotations of faithfulness judgments to create meta-evaluation benchmarks that measure the effectiveness of faithfulness metrics Fabbri et al. (2021); Pu et al. (2023); Tang et al. (2023); Goyal et al. (2023); Zhang et al. (2024c); Liu et al. (2023b, 2024b). This effort has led to the development of strong faithfulness metrics Laban et al. (2022); Fabbri et al. (2022); Zha et al. (2023). More recently, LLM-based metrics, which prompt powerful LLMs to perform evaluations Liu et al. (2023a); Wang et al. (2023); Fu et al. (2024), have shown high correlations with human judgments. Subsequently, MiniCheck Tang et al. (2024b) has focused on distilling such knowledge into smaller NLI models, combining both strong performance and efficiency. While most of these studies concentrate on standard summarization tasks, few address long-form summarization Krishna et al. (2023); Zhang et al. (2024b); Huang et al. (2024). Specifically, Zhang et al. (2024b) collate a meta-evaluation benchmark and develop an automatic metric. We further extend their work by incorporating additional benchmarks, analyzing the performance of the latest faithfulness metrics, and, more importantly, investigating sensitivity to positional bias.